OpenStudy (anonymous):

I am in need of help, please...... 2/7 = z/-3 (solve using the addition principle) ??

OpenStudy (saifoo.khan):

it can be solved by another method as well.

OpenStudy (saifoo.khan):

\[\frac27 = \frac{z}{-3}\]

OpenStudy (anonymous):

hello

OpenStudy (saifoo.khan):

Cross multiply, \[z \times 7 = 2 \times -3\] \[7z = -6\] z = -6/7

OpenStudy (saifoo.khan):

hi

OpenStudy (anonymous):

okay that will work as it wants it to be done in the addition principal?

OpenStudy (anonymous):

okay, I can understand that but why in the heck do they confuse you when asking that it be done solving using the addition principle?

OpenStudy (anonymous):

@saifo.khan

OpenStudy (saifoo.khan):

umm, let me see.

OpenStudy (anonymous):

hey I have two more problems can you help me with them?

OpenStudy (anonymous):

5x - x = x + 3x

OpenStudy (saifoo.khan):

sure.

OpenStudy (saifoo.khan):

5x - x = x + 3x 4x = 4x

OpenStudy (anonymous):

does it go any more to reduce? or just stays 4x=4x

OpenStudy (saifoo.khan):

idk in which form u need the answer.

OpenStudy (anonymous):

last one is, Q = p-q/2, for p

OpenStudy (anonymous):

sorry long lag on my end on the computer

OpenStudy (saifoo.khan):

2Q = p -q p = 2Q +q

OpenStudy (saifoo.khan):

no problem, this website has old tradition of lagging.

OpenStudy (anonymous):

this is a relativity formula from physics?

OpenStudy (saifoo.khan):

ummm, i dont remember. LOL

OpenStudy (anonymous):

oh okay, thanks

OpenStudy (saifoo.khan):

uw/

OpenStudy (anonymous):

lol, this is how it is written in the book, as follows

OpenStudy (anonymous):

Q = p-q over 2, for p

OpenStudy (saifoo.khan):

\[Q=p-\frac q2\]

OpenStudy (saifoo.khan):

\[p = Q + \frac q2\]

OpenStudy (anonymous):

so, is that how I would re-write it?

OpenStudy (saifoo.khan):

\[p = Q + \frac q2\]

OpenStudy (anonymous):

thanks so much for your help, hope to see you around again in here....have a good nite!!

OpenStudy (saifoo.khan):

you welcome!! Yeah, sure! ^_^

OpenStudy (saifoo.khan):

well it's morning here. LOL

OpenStudy (anonymous):

lol, hey are you still there???